OpenStudy (o-b-server):

Linear equation and graphing questions! Fan and Medal for right answer! Please help!

OpenStudy (o-b-server):

Write an equation for the line that is parallel to y = –5x + 5 and passes through (–6, 3). A. y = 5x + 27 B. y = 5x – 9 C. y = –5x – 27 D. y = –5x + 9 I don't have time for explanation I need an answer!

OpenStudy (animeasgardian):

The answer is C

OpenStudy (animeasgardian):

@O-B-Server

OpenStudy (o-b-server):

Do you think you could help me with another question?

OpenStudy (animeasgardian):

Yes

OpenStudy (o-b-server):

Alright give me a second to type it out

OpenStudy (o-b-server):

Theresa started making homemade cards to send to friends and family and to sell at the local craft fair. The scatter plot shows how many cards Theresa made each hour she worked. Find an equation of a reasonable trend line for the scatter plot. A. y + 14.9 = 2.9(x – 5) B. y – 5 = 2.9(x – 14.9) C. x + 14.9 = 2.9(y – 5) D. y – 14.9 = 2.9(x – 5) The attachment is the scatter graph.

OpenStudy (o-b-server):

And these if you can! Write an equation for the line through (–3, –5) with slope m = –1 over 3. A. y + 5 = x + 3 B. y – 5 = –1 over 3(x + 3) C. y + 5 = –1 over 3(x + 3) D. y + 5 = 1 over 3(x – 3) Write y = –4 over 5x + 7 in standard form using integers. A. 4x + 5y = 35 B. –5y = 4x + 35 C. –5x = 4y + 35 D. x = –4 over 5y + 7 Write y = –4 over 7x + 2 in standard form using integers. A. –7y = 4x + 14 B. x = –4 over 7y + 2 C. 4x + 7y = 14 D. –7x = 4y + 14

OpenStudy (animeasgardian):

you'll have to rewrite those, they came out weird

OpenStudy (o-b-server):

Refresh the page it'll fix it

OpenStudy (animeasgardian):

@Ethanxiaorox could you please help him, I have to go

OpenStudy (one098):

Make a new post for each new question.